OpenStudy (anonymous):

ATTENTION!!!!!!!:):)::):) WILL GIVE MEDAL PLUS FAN Determine if triangle DEF with coordinates D (2, 1), E (3, 5), and F (6, 2) is an equilateral triangle. Use evidence to support your claim. If it is not an equilateral triangle, what changes can be made to make it equilateral? Be specific.

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@whpalmer4 @agent0smith

OpenStudy (whpalmer4):

Find the length of each side using the formula for the distance between two points \((x_1,y_1)\) and \((x_2,y_2)\): \[D = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\] An equilateral triangle by definition has 3 equal-length sides.

OpenStudy (anonymous):

so for the first points How would it look when I plugg them in

OpenStudy (whpalmer4):

from D (2,1) to E (3,5) call D \((x_1,y_1)\) and E \((x_2,y_2)\) distance from D to E is \(\sqrt{(3-2)^2+(5-1)^2} = \sqrt{1^2+4^2} = \sqrt{17}\)

OpenStudy (anonymous):

ok can you help me with the distance for the next points please

OpenStudy (whpalmer4):

it's the same procedure. pick one point, use it as \((x_1,y_1)\). the other point is \((x_2,y_2)\).

OpenStudy (anonymous):

yes can you please write it out for me